The sides of the triangle ABC represents 3 velocities AC(magnitude 3.6m/s) corresponds to the velocity of a boat as observed by a stationary tourist on a bridge. BC (magnitude 1.5m/s) corresponds to the velocity of a cyclist crossing the bridge, once again as seen by the tourist. Calculate the magnitude of the velocity vector AB, the velocity of the boat relative to the cyclist.
1
Expert's answer
2020-11-23T10:29:59-0500

Given,

Veocity "AC=3.6m\/s" ,"BC=1.5m\/s,"


As for the tourist standing on the bridge

the traingle formed is right angle triangle in which "\angle B=90^{\circ}"


Using pythagoras theorem in "\Delta ABC"

"\Rightarrow AC^2=AB^2+BC^2\\\Rightarrow(3.6)^2=AB^2+(1.5)^2\\\Rightarrow AB^2=12.96-2.25=10.71\\\Rightarrow AB=\sqrt{10.71}=3.27m/s"


Hence the magnitude of velocity of "AB" is "3.27m\/s"